answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

How many face cards are in a deck?

12, plus 2 jokers.

There are 54 cards in all, 52 in the regular deck plus the 2 jokers.

Why was the venn diagram invented?

So that you could visualize the various permutations in boolean algebra.

An Urn contains 5 white and 7 black balls five balls drawn at randoma Find the distribution function of the probability distribution of no of white balls?

with replacement: binominal distribution f(k;n,p) = f(0;5,5/12) without replacement: hypergeometric distribution f(k;N,m,n) = f(0;12,5,5)

Frequency approach for probability?

Well, that's not much of a question. Perhaps you are asking: What is the frequency interpretation of probability? This is called the classical interpretation of probability. Given n independent and identical trials with m occurrences of of a particular outcome, then the probability of this outcome, is equal to the limit of m/n as n goes to infinity. If you are asking: How can probabilities be estimated given data, based on frequency approach? A table is constructed, with intervals, and the number of events in each interval is calculated. The number of events divided by the total number of data is the relative frequency and an estimate of probability for the particular interval.

What is the probability of having twins if your dad is a fraternal twin and you are a identical twin and your moms dad is a fraternal twin?

The probability is the same as if you had no history of twins in your family. There is no genetic component with identical twins, it is just something that happens. Fraternal twins occur if a mother "drops" two eggs in one month. Your father being a fraternal twin does not make you "drop" more than one egg in a month and so does not increase your chances of having twins. Your maternal grandfather being a fraternal twin did not effect your mother's ability to drop more than one egg and so did not effect your ability to "drop" more than one egg and so does not increase your ability to have twins.

What does it mean for a sample to have a standard deviation of zero?

It means that there are is no variation from the mean. In other words, all values in your sample are identical.

What is the probability of getting 2 girls and 2 boys of 4 children?

It depends. If the order is important, (2 girls then 2 boys): There's a 1/2 chance of any sex, so the probability is 1/16(or 6.25%). But if it's just any order of 2 girls and 2 boys, then it's a different outcome. Consider the following table. The desired outcome is 2 girls, 2 boys in any order. We can just look for 2 girls, and convert to Boolean values of 1 for girl and 0 for boy. Here is the table:

0 0 0 0: 0

0 0 0 1: 1

0 0 1 0: 1

0 0 1 1: 2*

0 1 0 0: 1

0 1 0 1: 2*

0 1 1 0: 2*

0 1 1 1: 3

1 0 0 0: 1

1 0 0 1: 2*

1 0 1 0: 2*

1 0 1 1: 3

1 1 0 0: 2*

1 1 0 1: 3

1 1 1 0: 3

1 1 1 1: 4

Note each row has five numbers. The first four numbers represent the four children, and the fifth number is a sum of these four numbers. Any time the sum equals exactly 2, then the desired outcome (2 out of the four children are girls).

This happens 6 times out of the 16 total possible outcomes, so the probability is 6/16, or 3/8 = 37.5%

What is the probability of drawing 2 kings from a standard deck of fifty-two cards without replacement?

Combination of 52 things taken 2 at a time is (52 x 51)/(2 x 1)=1326. This is the total possible outcomes when drawing two cards from a deck. Combination of 4 things taken 2 at a time is (4 x 3)/(2 x 1) = 6. This is the total possible combinations of two kings (Heart/Diamond, Heart/Spade, Heart/Club, Diamond/Spade, Diamond/Club, Spade/Club). The probability, then, is 6 out of 1326, or 1/221.

What is the probability of having 3 daughters in a row?

it one out 8 1/8 bc

first child 1/2

second child 1/2

third child 1/2

1/2 x 1/2 x 1/2 = 1/8

What is the difference between a probability distribution and sampling distribution?

A sampling distribution function is a probability distribution function. Wikipedia gives this definition: In statistics, a sampling distribution is the probability distribution, under repeated sampling of the population, of a given statistic (a numerical quantity calculated from the data values in a sample). I would add that the sampling distribution is the theoretical pdf that would ultimately result under infinite repeated sampling. A sample is a limited set of values drawn from a population. Suppose I take 5 numbers from a population whose values are described by a pdf, and calculate their average (mean value). Now if I did this many times (let's say a million times, close enough to infinity) , I would have a relative frequency plot of the mean value which will be very close to the theoretical sampling pdf.

You randomly select a month of the year What is the probability that you select a month with 28 or 30 days in it?

First we don't consider leap year

There is 1 month with 28 days: february

There are 4 months with 30 days: april, june, september, november

So the probability to select a month with 28 or 30 days is (1+4)/12 = 5/12

If the year is a leap year then the probability is 4/12

How do you solve this problem there are 3 sets of examination in a vocational TESDA course the probability that an examinee will pass in each subject is 80 what is the probability that an examinee pas?

The problem as I understand it is the examinee must pass test 1 and test 2 and test 3. Therefore, the probability of passing all 3 exams is .8*.8*.8 = 0.512.

What is the probability of P z-0.41 using the standard normal distribution?

See the related link for the area at 0.41 (same as -0.41) which is 0.1591. This area, which is the probability, is from minus infinity to -0.41. If you want the area from -0.41 to plus infinity you need to take 1 - 0.1591 which is 0.8409.

Trending Questions
What is the probability of how many ways are there of choosing an even number from the first 10 positive whole numbers? Why are standard z values so important? What is the difference between population distribution and sample distribution an d sampling distribution? Which of the following is the best definition of sample space of a probability event? If youu had to cancel a picnic because of rain what has changed your plans? Probability 1 of 90 percent? What z-score separates the top 10 percent from the rest of the scores? Can you have a period 4 days after conception? If a bag contains 3 green marbles 5 blue marbles and 2 yellow marbles find the probability that the marble is blue? 7 is to 34 as 5 is to? How many Non spades are in deck of cards? How two same ellipses have rolling contact? What is the probability 1 out of 8? Each of the numbers below is a solution of the inequality 2x plus 3 greater than 7 except A 3 B 6 C 1 D 5? What is the probability of randomly guessing at least 8 questions correct on a 10 question true-or-false quiz? WHAT IS SAMPLE SPACE WHEN A COIN IS TOSSED ONCE? What would be the color vision of children if their mom is colorblind and their dad is normal? How many 5 numbers combinations can you make from the numbers 1 to 49 and you can use multiple combinations but still use each number only once? How do firearms function? If a test has a normal distribution with a mean of 400 and a standard deviation of 100 what would be the minimum score for the top 3 percent of students?